You are thinking of building a new machine that will save you $2,000 in the first year. The machine will then begin to wear out so that the savings decline at a rate of 3% per year forever. What is the present value of the savings if the interest rate is 9% per year? The present value of the savings is s(Round to the nearest dollar.)
Answer #1

Value after first year=(Savings in first year(1+Growth rate))/(Interest Rate-Growth Rate)

=(2000(1-0.03))/(0.09-(0.03))

=(1940/0.12)

=$16,166.67(Approx)

Hence present value of savings=Future value*Present value of discounting factor(rate%,time period)

=2000/1.09+$16,166.67/1.09

which is equal to

=$16,667(Approx).